If interest on a $4,500 loan is $720 what would be the interest on a loan of $ 5, 800
Answer by Theo(13342)

720/4500 = x/5800

cross multiply to get 720 * 5800 = 4500 * x

solve for x to get x = (720 * 5800) / 4500 = 928

assuming the interest rate is the same, then the interest on 5800 should be 928.

let's see if that's true.

interest rate on 720/4500 is equal to .16

.16 * 5800 is equal to 928.

it works.

this is a direct variation type equation.

y = k*x

when y is 720 and x is 4500, then:

720 = k * 4500.

solve for k to get k = 720 / 4500 = .16

now apply k to your new value for x.

y = k * x becomes y = .16 * 5800 which becomes y = 928.